OpenStudy (anonymous):

The function f(x) = 9x+5/x-8 is one to one find an equation for f^-1(x), the inverse function

OpenStudy (asnaseer):

first write this as:\[y=\frac{9x+5}{x-8}\]then swap the x and y's to get:\[x=\frac{9y+5}{y-8}\]now just rearrange this to get it in the form: y = f(x)

OpenStudy (asnaseer):

that will be your inverse function.

OpenStudy (anonymous):

im having trouble simplifying x = 9y+5 ----- y-8 how do i solate y?

OpenStudy (anonymous):

@dylanlamoreaux , multiply both sides by (y-8) then distribute the x into y-8 on the left side

OpenStudy (campbell_st):

\[x = \frac{9y + 5}{x -8}\] multiply both sides by (y - 8) \[x(y - 8) = 9y + 5\] expand and simplify xy - 8x = 9y + 5 y( x -9) = 8x + 5 \[y =\frac{8x + 5}{x - 9}\]

OpenStudy (asnaseer):

Actually what @campbell_st said earlier is true - I didn't read the question fully. This is NOT a one-to-one function as the original function f(x) = (9x+5)/(x-8) leads to f(8) = \(\pm\infty\) as x approaches 8 from the left/right.